If f(-4)=5, which of the following could not be the inverse of f(x)?
a)g(x)=x-9
b)g(x)=x^2-11
c)g(x)=x^2-5x-4
d)g(x)=-x^2+3x+6
Please help. My textbook doesn't have an example for a problem like this so I don't know where to even start.

Found 2 solutions by ikleyn, Edwin McCravy:
Answer by ikleyn(52802)

.
You need to check which of the listed functions has the output -4 when the input is 5.

If the function from the list HAS the output  -4,  then it potentially CAN BE the inverse to  f(x).

It the function from the list HAS NO the output  -4,  it CAN NOT be the inverse to f(x).     Then it is your lovely function.


Answer by Edwin McCravy(20060)

She is forgetting that only one-to-one functions can HAVE or BE 
inverse functions. These are functions  which pass both the 
horizontal and vertical line tests.  No 2nd degree polynomial 
functions are one-to-one, because they have u-shaped graphs 
(parabolas) which do not pass the horizontal line test (unless
their domain is restricted.

So b), c), and d) are not one-to-one functions, and are immediately
eliminated.  The only possibility is a).  But let's see why.

f(-4)=5 means this: 

When -4 is substituted for x in f(x), the result is 5.

Then in order for g(x) to be the inverse of f(x), then g(x)
must be a one-to-one function and we have the vice-versa of
the above.  That is:

when 5 is substituted for x in the right side of g(x), the result
must be -4.

a) g(x)=x-9
   g(5)=5-9 = -4

So (a) is the correct answer.

Notice that choices c) and d) might trick a student into thinking
they are correct because you also get -4 when you substitute x=5

c)g(5)=5^2-5(5)-4 = 25-25-4 = -4
d)g(5)=-5^2+3(5)+6 = -25+15+6 = -4

But they are not one-to-one functions because a u-shaped graph cannot 
pass the horizontal line test.

Edwin